You take a 5/1 interest-only ARM for $250,000, monthly payments, 30-year term. The initial contract rate is 4.5% for the first 5 years. What is your monthly payment for the first 5 years?
Answer #1

Mortgage Amount (P)= 250000

Year of Amortization = 30

Number of months (n)= 30*12= 360

Rate applicable for first 5 years will be Initial contract rate

=4.5%

Monthly rate (i)= 4.5%/12=0.00375

Monthly payment formula = P*i/(1-((1+i)^-n))
250000*0.00375/(1-((1+0.00375)^-360))

=1266.713275

So Payment for first 5 years will be $1266.71

Note: Excel or Financial calculator function =

=Pmt(rate per period, Number of periods, -Present Value)

Pmt(0.00375, 360, -250000)

=$1,266.71

(Please thumbsup)